An insurance agent receives a 12% fee on each new policy payment and a 3% fee on each renewal payment. If the agent received $1350 fee on total payments of $26250, how much were the payments for new policies and renewal payments?
Answer by mananth(16946)

Let $ x be the new policies
$26250 -x will be the renewals.
..
12% on new
3 % on renewal.
Amount received = $1350
..
0.12x+0.03(26250-x)= 1350
0.12x+787.5-0.03x = 1350
0.09x = 1350-787.5
0.09x = 562.5
divide by 0.09
x= $6,250 new policies
renewals will be 26250-6250=20,000